The weights of students in a certain school are normally distributed with a mean weight of 66 kg. 10% have a weight greater than 70kg. What percentage of students weighs between 62kg and 66kg?


1
Expert's answer
2022-04-16T04:13:04-0400

"P(X>70)=0.1,\\
P(X<70)=1-P(X>70)=\\=1-0.1=0.9."

From z-table the 90th percentile is z = 1.282.

So,

"z=\cfrac{x-\mu} {\sigma}, \\
1.282=\cfrac{70-66}{\sigma},\\
\sigma=\cfrac{4} {1.282} =3.12,\\
z_1=\cfrac{62-66} {3.12}=-1.28, \\
z_2=\cfrac{66-66}{3.12}=0,\\"

"P(62<X<66) =P(-1.28<Z<0)=\\
=P(Z<0)-P(Z<-1.28)=\\=0.5-0.1003=0.3997=39.97\%\\\text{ (from z-table). }"